OpenStudy (anonymous):

Given f(x)= -5x+3, find f(a+1)

OpenStudy (hba):

Welcome to OS, Just plug in a+1 in place of x :) f(x)=-5x+3 f(a+1)=

OpenStudy (anonymous):

I still don't get how to determine the value of a

OpenStudy (hba):

We don't have to determine the value of a

OpenStudy (anonymous):

so I just use f(a+1)=-5x+3?

OpenStudy (hba):

Yeah you just plug in a+1 in place x :)

OpenStudy (anonymous):

ok, thanks guys :)

OpenStudy (hba):

What did you get @Aairo

OpenStudy (anonymous):

f(a+1)=-5a+4

OpenStudy (hba):

Please show me your working.

OpenStudy (hba):

I also know that,He did not distribute correctly @Hero

OpenStudy (hba):

See a(b+c)=ab+ac